# COSMOS: Cross-Modality Self-Distillation for Vision Language Pre-training

### v1

@@ E.3 Experiment on PixelProse Dataset
To demonstrate the adaptability of COSMOS to various captions, we employ PixelProse [65], a synthetic caption dataset similar to DreamLIP. PixelProse filters and combines data from CommonPool [25], CC12M [64], and RedCaps [19] to create over 16 million image and alt-text pairs. Singla et al. [65] used Gemini 1.0 Pro [65] to generate new captions. As shown in Tab. 12 and Tab. 13, we trained CLIP [59], SigLIP [80], and COSMOS on PixelProse using the same setup as the previous experiments. The actual number of image-text pairs used to train these models is detailed in Tab. 6.
In Tab. 12, COSMOS consistently outperforms CLIP and SigLIP models by a significant margin with both ViT-B/16 and ViT-B/32 vision encoders. For instance, COSMOS with ViT-B/16 achieves 62.4% and 43.4% in R@1 scores for image-to-text and text-to-image retrieval on MSCOCO. This performance even surpasses strong baselines from OpenCLIP [13], which achieve (56.5% and 37.9%) on LAION-400M, (58.2% and 39.8%) on DataComp-1B, and (59.3% and 41.7%) on LAION-2B (see Tab. 1 main).
Similarly, Tab. 13 shows that COSMOS outperforms previous baselines in the zero-shot classification tasks. Notably, COSMOS achieves the best average accuracy of 60.7% with ViT-B/16 and 58.0% with ViT-B/32 when training on PixelProse, as opposed to training on the other long caption datasets proposed by DreamLIP [83] (CC3M, CC12M, YFCC15M, and Merged-30M). However, the same models trained on PixelProse do not achieve the best accuracy in retrieval tasks according to Tab. 1 main and Tab. 10. We infer that this may be due to the number and content of the captions used in training. DreamLIP utilizes three MLLMs to generate long synthetic captions, resulting in longer captions than those generated by PixelProse, which only uses Gemini Pro. Consequently, models trained on longer and more varied captions are better at capturing local information, while models trained on relatively shorter and similar captions are better at focusing on global objects in the image. Investigating this trade-off based on the nature of these long captions would be an intriguing direction for future work.
@@
Table 11: Zero-shot classification results in terms of top-1 accuracy on ImageNet [18] and Food101 [4], CIFAR-10 [38], CIFAR-100 [38], SUN397 [73], Stanford Cars [37], FGVCAircraft [50], DTD [14], Oxford Pets [57], Caltech101 [24], Flowers102 [54], and ImageNet [18] datasets. The vision encoder architecture is ViT-B/32. The best results are highlighted in bold. Results are reproduced with our setup for fair comparison unless otherwise marked. : Results copied from their work. : Results obtained using their official pre-trained weights.
@@
2 globals + 0 locals 31.0 50.1 35.4 80.9 63.2 14.6h 18.2G
2 globals + 2 locals 34.8 51.5 38.5 82.7 65.9 21.4h 23.4G
2 globals + 4 locals 36.3 52.3 39.6 83.5 67.4 26.5h 28.2G
2 globals + 6 locals 37.1 53.1 40.1 84.1 68.6 31.7h 32.6G
Table 14: Ablation on different numbers of local crops. We show total running time and peak memory usage per GPU using ViT-B/16. We train our model on four 4-GPU machines using the CC3M dataset with a batch size of 1,024.
@@
Table 15: Evaluation on Winoground [68]
@@
Table 16: Ablation on loss scale of SILC [53] . SILC with different loss scale (a,b) where the total loss is calculated as $\mathcal{L}_{\text{total}}=a\mathcal{L}_{\text{CLIP}}+b\mathcal{L}_{\text{self-distill}}$. Models are trained on CC3M with batch size of 1,024 and 1 global and 1 local crops are used as augmentation.
E.4 Experiment on Winoground Dataset
In Tab. 15, we compare COSMOS with three baselines (CLIP [59], SigLIP [80], and DreamLIP [83]) as well as OpenCLIP [13] models trained on larger datasets. The results for OpenCLIP models indicate that scaling up the size of pre-training datasets does not necessarily improve the performance. In other words, large datasets alone cannot ensure high compositional knowledge in VLMs. However, COSMOS outperforms all OpenCLIP models, suggesting that our self-distillation and cross-modality learning effectively enhance the model’s contextual understanding. COSMOS achieves 30.8% and 16.5% in text and image scores, respectively, surpassing the previous best baselines of 29.3% text score from CLIP and 15.8% image score from DreamLIP, while achieving comparable results in terms of group score.
E.5 Ablation Study on Locap Crops
In Tab. 14, we report the GPU memory requirements and training time based on the number of local crops, along with the corresponding zero-shot performance. We fixed the number of global crops for both image and text at 2, while varying the number of local crops included during training. As a result, increasing the number of local crops improves zero-shot classification and retrieval tasks, achieving 37.1% accuracy on ImageNet and 53.1% and 40.1% R@1 scores on image-to-text and text-to-image retrieval on the MSCOCO validation set. The improvement is most significant between 0 and 2 local crops, with diminishing returns as more local crops are added. GPU memory usage and training time also increase with the number of local crops. Therefore, one should determine the optimal number of local crops based on available computational resources.
E.6 Rescaling Loss in Previous Methods
@@
Similarly, other works [43, 52, 21, 61] that integrated self-supervised learning in VLM training adopt loss scaling parameters to balance the learning speed between the contrastive objective and the self-distillation objective. This is due to the different scales of the loss functions, as previous works often employ symmetric InfoNCE loss [56] for contrastive learning while using cross-entropy loss between masked inputs or global and local crops for self-supervised learning. We unify the loss function with the InfoNCE loss, eliminating the need for a scaling factor, since the CLIP loss and the COSMOS loss are already updated on the same scale.

### v2

@@ E.3 Experiment on PixelProse Dataset
To demonstrate the adaptability of COSMOS to various captions, we employ PixelProse [70], a synthetic caption dataset similar to DreamLIP. PixelProse filters and combines data from CommonPool [25], CC12M [68], and RedCaps [19] to create over 16 million image and alt-text pairs. They used Gemini 1.0 Pro [70] to generate new captions. As shown in Tab. 12 and Tab. 13, we trained CLIP [63], SigLIP [88], and COSMOS on PixelProse using the same setup as before. The actual number of image-text pairs used to train these models is detailed in Tab. 6.
In Tab. 12, COSMOS consistently outperforms the CLIP and SigLIP models by a significant margin with both ViT-B/16 and ViT-B/32 vision encoders. For instance, COSMOS with ViT-B/16 achieves 62.4% and 43.4% in R@1 scores for image-to-text and text-to-image retrieval on MSCOCO. This performance even surpasses strong baselines from OpenCLIP [13], which achieve 56.5% and 37.9% on LAION-400M, 58.2% and 39.8% on DataComp-1B, and 59.3% and 41.7% on LAION-2B (see Tab. 1 main).
Similarly, Tab. 13 shows that COSMOS outperforms previous baselines in the zero-shot classification tasks. Notably, COSMOS trained on PixelProse achieves the best average accuracy of 60.7% with ViT-B/16 and 58.0% with ViT-B/32 among all models trained on the synthetic caption datasets (e.g., CC3M, CC12M, YFCC15M, and Merged-30M). However, the same models trained on PixelProse do not achieve the best accuracy in retrieval tasks according to Tab. 1 main and Tab. 10. We infer that this might be due to the length and content of the captions used during the training. DreamLIP utilizes three MLLMs to generate long synthetic captions, resulting in longer captions than those generated by PixelProse, which only uses Gemini Pro. Consequently, models trained on longer and more varied captions are better at capturing local information (i.e., better at retrieval task), while models trained on relatively shorter and similar captions are better at focusing on global objects in the image (i.e., better at classification task). Investigating this trade-off based on the nature of these long captions would be an intriguing direction for future work.
@@
Table 11: Zero-shot classification results in terms of top-1 accuracy on the ImageNet [18] and Food101 [4], CIFAR-10 [39], CIFAR-100 [39], SUN397 [79], Stanford Cars [38], FGVCAircraft [54], DTD [14], Oxford Pets [61], Caltech101 [24], Flowers102 [58], and ImageNet [18] datasets. The vision encoder architecture is ViT-B/32. The best results are highlighted in bold. Results are reproduced with our setup for fair comparison unless otherwise marked. : Results copied from their work. : Results obtained using their official pre-trained weights.
@@
1 CLIP: 1 globals + 0 locals 24.5 38.9 26.8 69.1 51.2 3.8h 10.5G
2 CLIP: 2 globals + 0 locals 29.7 47.6 33.8 78.1 60.7 12.2h 16.3G
3 CLIP: 2 globals + 2 locals 31.3 49.6 35.9 78.4 62.1 17.6h 20.6G
4 COSMOS: 2 globals + 0 locals 31.0 50.1 35.4 80.9 63.2 14.6h 18.2G
5 COSMOS: 2 globals + 2 locals 34.8 51.5 38.5 82.7 65.9 21.4h 23.4G
6 COSMOS: 2 globals + 4 locals 36.3 52.3 39.6 83.5 67.4 26.5h 28.2G
7 COSMOS: 2 globals + 6 locals 37.1 53.1 40.1 84.1 68.6 31.7h 32.6G
Table 14: Ablation on the training efficiency of COSMOS. Total running time and peak memory usage per GPU are shown with different configuration of global and locap crops. We also report zero-shot classification results on ImageNet [18] and zero-shot retrieval results on Flickr30K [84] and MSCOCO [48]. We train our model (ViT-B/16) on four 4-GPU machines using the CC3M dataset with the batch size of 1,024.
E.4 Ablation Study on Training Efficiency
In Tab. 14, we report the GPU memory requirements and training time based on the number of global and local crops, along with the corresponding zero-shot performance. Compared to CLIP, COSMOS requires 10-20% more training time and GPU memory due to its teacher-student setup (row 2 vs row 4, and row 3 vs row 5). As no gradients flow into the teacher during training, COSMOS does not add too much computational overhead while outperforming CLIP.
For COSMOS, we fixed the number of global crops for both image and text to two, while varying the number of local crops included during training (row 4-row 7). As a result, increasing the number of local crops improves zero-shot classification and retrieval tasks, achieving 37.1% accuracy on ImageNet and 53.1% and 40.1% R@1 scores on image-to-text and text-to-image retrieval on the MSCOCO validation set with six local crops. The improvement is most significant between zero and two local crops (row 4 vs row 5), with diminishing returns as more local crops are added. GPU memory usage and training time also increase with the number of local crops. Therefore, one could determine the optimal number of local crops based on the available computational resources.
E.5 Experiment on the Winoground Dataset
@@
Table 15: Evaluation on Winoground [73].
In Tab. 15, we compare COSMOS with three baselines (CLIP [63], SigLIP [88], and DreamLIP [91]) as well as OpenCLIP [13] models trained on larger datasets. The results of the OpenCLIP models indicate that scaling up the size of pre-training datasets does not necessarily improve the performance. In other words, large datasets alone cannot ensure high compositional knowledge in VLMs. However, COSMOS outperforms all OpenCLIP models, suggesting that our self-distillation and cross-modality learning effectively enhance the model’s contextual understanding. COSMOS achieves 30.8% and 16.5% in text and image scores, respectively, surpassing the previous best baselines of 29.3% text score from CLIP and 15.8% image score from DreamLIP, while achieving comparable results in terms of group score.
E.6 Experiment on MLLM setting
Method Data Size ScienceQA POPE GQA TextVQA MMMU
CLIP [63] 30M 65.2 80.1 58.7 53.5 33.9
OpenCLIP [13] 400M 67.2 81.3 59.8 54.4 36.5
COSMOS 30M 67.8 83.2 60.4 55.3 36.8
Table 16: LLaVA v1.5 experiment. CLIP and COSMOS trained on Merged-30M and OpenCLIP trained on LAION-400M.
In order to evaluate COSMOS on MLLM setting, we adapt our vision encoder (ViT-B/16) to LLaVA framework [49]. We follow the training process of LLaVA v1.5 which consists of feature alignment pre-training and visual instruction tuning. Various benchmarks are selected for comprehensive evaluation including ScienceQA [53], POPE [46], GQA [34], TextVQA [69], and MMMU [86]. In Tab. 16, COSMOS outperforms both CLIP trained on the same data (Merged-30M) and OpenCLIP trained on LAION-400M, which demonstrates the effectiveness of COSMOS on visual reasoning and compositional question answering.
E.7 Ablation Study on Text Cropping Strategy
Method ImageNet MSCOCO Flickr30K
Top-1 I2T@1 T2I@1 I2T@1 T2I@1
Masked text [45] 28.5 46.0 32.8 76.2 59.2
Summarized text [26] 33.9 51.7 37.8 81.0 65.3
Local within global 34.6 52.1 37.8 81.2 66.1
COSMOS 35.1 52.6 38.9 83.0 66.5
Table 17: Ablation on text cropping strategies. We compare various text cropping methods in terms of zero-shot classification results on ImageNet [18] and zero-shot retrieval results on Flickr30K [84] and MSCOCO [48]. All models are trained on CC3M with batch size 1,024 and ViT-B/16 image encoder.
In our paper, both global and local crops of captions are randomly sampled from long synthetic captions. Typically, global texts (1-5 sentences) are longer than local texts (1 sentence). The sampling processes of global and local crops are entirely independent. Therefore, global captions may or may not include local captions, similar to the image crop method in SimCLR [11] and DINO [7]. While mismatches can occur between global and local captions via random sampling, as with the global and local crops of images, the model can learn conceptual correspondences. For example, for an image of a park, the global text may describe the park, while the local text describes a dog. The model then learns that "park" and "dog" are related in the text.
[…]
E.8 Additional Results on Semantic Segmentation
Method VOC20 City. Context59 ADE COCO-Stf.
CLIP [63] 11.3 5.0 4.5 1.3 2.8
SigLIP [88] 14.5 5.5 5.8 2.2 3.8
DreamLIP [91] 1.8 0.9 0.4 0.1 0.1
COSMOS 53.6 13.9 15.7 8.5 10.7
Table 18: Zero-shot semantic segmentation results in terms of mean Intersection over Union (mIoU). The vision encoder architecture is ViT-B/16 and all models are trained on Merged-30M.
In addition to Table 3 in main paper, we compare segmentation performance of COSMOS to CLIP [63], SigLIP [88], and DreamLIP [91] in Tab. 18. Surprisingly, DreamLIP performs poorly in every semantic segmentation benchmarks, likely because its loss function weakens local image representation by matching local visual patches with global text. Ours is by far the best performing method, likely due to its cross-modality embedding and local-to-global matching, alleviating local feature suppression.
E.9 Additional SOTA Comparison
Method Data Size Batch Size MSCOCO Flickr30K
I2T T2I I2T T2I
VeCLIP [40] 300M 32k 67.8 48.9 91.2 76.3
MobileCLIP-B [76] 1B 65k 68.8 50.6 91.4 77.3
COSMOS 30M 4k 68.0 52.5 92.9 80.3
Table 19: Comparison to VeCLIP [40] and MobileCLIP [76] in terms of zero-shot retrieval results on Flickr30K [84] and MSCOCO [48].
In Table 1 and 2 in the main paper, COSMOS is compared to other methods using multi-modal data augmentation including DreamLIP [91], LaCLIP [23], and MLLM-A [51]. We additionally report the results of VeCLIP [40] and MobileCLIP [76] for comparison. VeCLIP exploits LLaVA [50] to generate detailed captions while MobileCLIP utilizes CoCa [85] to generate multiple synthetic cations. Although COSMOS is trained with much smaller pre-training set and batch size, it outperforms other methods, demonstrating the efficient usage of synthetic captions within our framework.
E.10 Rescaling Loss in Previous Methods
@@
Table 20: Ablation on the loss scale of SILC [57]. SILC with different loss scale (a,b) where the total loss is calculated as $\mathcal{L}_{\text{total}}=a\mathcal{L}_{\text{CLIP}}+b\mathcal{L}_{\text{self-distill}}$. Models are trained on CC3M with the batch size of 1,024 and one global and one local crops are used as augmentation.
@@
Similarly, other works [45, 56, 21, 65] that integrated self-supervised learning in VLM training adopt loss scaling parameters to balance the learning speed between the contrastive objective and the self-distillation objective. This is due to the different scales of the loss functions, as previous works often employ the symmetric InfoNCE loss [60] for contrastive learning while using the cross-entropy loss between masked inputs or global and local crops for self-supervised learning. We unify the loss function with the InfoNCE loss, eliminating the need for a scaling factor, since the CLIP loss and the COSMOS loss are already updated on the same scale.